OpenStudy (anonymous):

Help please √x - 2 + 5 = 6

OpenStudy (anonymous):

x=3

OpenStudy (anonymous):

again please use parenthesis to clarify the part of the equation the square root encompasses

OpenStudy (anonymous):

I DONT KNOW HOW

OpenStudy (anonymous):

is x-2 inside the sqrt?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

this is what i meant √(x - 2) + 5 = 6

OpenStudy (anonymous):

what are the steps to solving

OpenStudy (anonymous):

the goal is to solve for x this means isolate x so its alone on one side this can be done by adding/ subtracting/ multiplying/ dividing /etc things to BOTH sides now the first step √(x - 2) + 5 = 6 we want to move the 5 on the left side over to the right side this can be done if we subtract 5 to both sides of the equation what happens is on the left side, 5-5 =0 so it cancels out on the left side on the right side, we get 6-5 √(x - 2) + 5-5 = 6-5 √(x - 2) = 1

OpenStudy (anonymous):

the next step would be to get rid of that square root any ideas?

OpenStudy (anonymous):

subtract

OpenStudy (anonymous):

actually, you want to square both sides take both sides to the 2nd power

OpenStudy (anonymous):

i did the rest:)

OpenStudy (anonymous):

\[(\sqrt{x} )^2 = x\]

OpenStudy (anonymous):

anyways once you square it you get x-2 =1 then you add 2 to both sides which cancels out the -2 on the right side to get x=3